OpenStudy (lasttccasey):

Calc Question: I have this equation, -ln|100-y| = (2x^(3/2))/3+C and I need to get solve for y but I'm not sure how to get rid of the -ln without putting it in a fraction like 1/(100-y) by taking the exponential root.

OpenStudy (lasttccasey):

\[-\ln|100-y| = \frac{ 2x ^{3/2} }{ 3 }+C\]

OpenStudy (anonymous):

Let me rephrase that. What section of Calc I are you in right now.

OpenStudy (anonymous):

I.e. what chapter/topic.